The ratio of two numbers is 7 : 2 the difference between the two numbers is 15 what is the greater number
Answer by mananth(16946)

let the numbers be x & y
x/y = 7/2
x=7y/2
x-y =15
plug x
7y/2 - y=15
multiply by 2
7y-2y=30
5y=30
/5
y=6
Now
x-y=15
x-6=15
x=15+6
x=21